OpenStudy (anonymous):

More with functions: QUICK QUESTION Which of the equations do I do first? G or F? For example: If f(x)=x+4, g(x)=5x-1 what is f + g of -3? I know how to do the equation but I don't know which one I should substitute -3 for x

OpenStudy (anonymous):

\[(f+g)(-3)=f(-3)+g(-3)\]

OpenStudy (anonymous):

so f first right?

OpenStudy (anonymous):

is it always going to be f first?

OpenStudy (anonymous):

because one of the equations I tried said what is g of f of -1 and I had to do F first

OpenStudy (anonymous):

guys please help :(

OpenStudy (jdoe0001):

hmm, is an addition, the order in whiich the sum will be conducted won't matter

OpenStudy (austinl):

It really doesn't matter which one you do first for this particular problem.

OpenStudy (jdoe0001):

2+ 3 = 5 3 + 2 = 5 x + 3x = 4x 3x + x = 4x

OpenStudy (anonymous):

okay if its addition it doesn't matter what if its multiplication?

OpenStudy (jdoe0001):

on multiplication, it won't matter, if you end up doing f(x) * g(x) same, the order doesn't matter in multiplication

OpenStudy (anonymous):

but I tried one that said f of g of 2 im guessing it was multiplication and i did it both ways and I got two different answers

OpenStudy (jdoe0001):

hehe, w00t, same 2 functions?

OpenStudy (austinl):

\(f(g(2))\)?

OpenStudy (austinl):

You first calculate g(2), and then you plug that value in for x in f(x).

OpenStudy (anonymous):

hmm yeah like that austin

OpenStudy (jdoe0001):

ohhh, that's not multiplication then

OpenStudy (anonymous):

yeah thats what I did but how did you know you do g first?

OpenStudy (anonymous):

oohhhh is it because its in parantheses?

OpenStudy (jdoe0001):

because is g(2) is the DOMAIN of f(x), so you evaluate g(2) first, so you can pass it to f(x)

OpenStudy (anonymous):

what do you mean domain?

OpenStudy (austinl):

You have \(f(x)\) and \(g(x)\) If you have \(f(g(x))\), that means that you would plug in \(g(x)\) for all instances of x in \(f(x)\). BUT because in your problem, it is \(g(2)\), that means that you have to plug in 2 for all instances of x in \(g(x)\). Then you plug that value into \(f(x)\).

OpenStudy (anonymous):

I get that plug in part and how to get the answer I just don't know which one to do first wether its f first or g first

OpenStudy (jdoe0001):

\(\bf f(x) = x^4 + x^3 + x^2...\\ g(x) = whatever\\ f(g(x)) = (whatever)^4 + (whatever)^3 + (whatever)^2 ...\)

OpenStudy (austinl):

Find \(f(g(3))\) \(f(x)=2x+3\) \(g(x)=5x-2\) \(g(3)=5(3)-2=13\) \(f(g(3))=2(13)+3=29\)

OpenStudy (anonymous):

Ii understand how to do that guys its just I don't know which to plug in FIRST

OpenStudy (anonymous):

okay lets say it said g(f(3) I would plug in 3 into f first right?

OpenStudy (austinl):

If it is \(f(g(\text{insert number here}))\) You calculate \(g(\text{insert number here})\) first, then you plug that into the greater value.

OpenStudy (anonymous):

ohh okay I think I get it...was i right about what I said before?

OpenStudy (austinl):

Mindless__Unique: okay lets say it said g(f(3) I would plug in 3 into f first right? This?

OpenStudy (anonymous):

yes

OpenStudy (austinl):

Yes, you would plug 3 into f(x) first.

OpenStudy (anonymous):

Okay I got it now :D THANKS SO MUCH YOU GUYS YOU SAVED MY retrice

OpenStudy (austinl):

Watch the language lol.

OpenStudy (anonymous):

How do I give you guys your medals?

OpenStudy (anonymous):

lol opps

OpenStudy (austinl):

well, you can only give one. I can give one to whoever.

OpenStudy (anonymous):

Thanks again guys

OpenStudy (austinl):

No problemo.